HUBBLE SPACE TELESCOPE DAILY REPORT #2529 PERIOD COVERED: 0000Z (UTC) 12/29/99 - 0000Z (UTC) 12/30/99 Daily Status Report as of 364/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8491 (Transition to hold, Decon and Cool Down Procedure) This program covers the transition from Protect Safe Mode to Hold Mode, then performs a Protect Decontamination. This is followed by a series of UV monitors and focus checks at -43C, then a special decon to take the instrument to -88C. The program completed nominally. 1.2 Completed STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.3 Completed Three Sets of WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 1.4 Completed STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.5 Completed Three Sets of STIS/MA1/MA2 8504 (SM3A MAMA Turn-on) The Space Telescope Imaging Spectrograph (MA1/MA2) was used to turn on MAMA high voltages and take darks three times per day. The observations completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 24 Successful: 24 2.3 Operations Notes: 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW